OpenStudy (anonymous):

Nicole is on her way home in her car. She has driven 20 miles so far, which is four-fifths of the way home. What is the total length of her drive?

OpenStudy (anonymous):

4/5 = 20 miles 1/5 = 20/4 = 5 miles 5/5 = 5miles x 5 = 25 miles.(total length)